Inventory information for Part 311 of Bonita Corp. discloses the following information for the month of June. June 1 Balance 304 units @ $13 June 10 Sold 205 units @ $30 11 Purchased 804 units @ $15 15 Sold 504 units @ $32 20 Purchased 498 units @ $16 27 Sold 295 units @ $34 (a) Assuming that the periodic inventory method is used, compute the cost of goods sold and ending inventory under (1) LIFO and (2) FIFO. (1) LIFO (2) FIFO Cost of Goods Sold $enter a dollar amount $enter a dollar amount Ending Inventory $enter a dollar amount $enter a dollar amount
Answer #1
Unit Unit Cost Total Cost
June 1 304 13 3952
June 11 804 15 12060
June 20 498 16 7968
Total 1606 23980

Sale unit= 205+504+295 = 1004

Calculate following

LIFO FIFO
Cost of goods sold (506*15+7968) = 15558 (700*15+3952) = 14452
Ending inventory 23980-15558 = 8422 23980-14452 = 9528